OpenStudy (josefeth00):

the sum of a number and its reciprocal is 61/30. find the number

OpenStudy (mathstudent55):

Start by choosing a variable. Do you like x?

OpenStudy (josefeth00):

yea I wrote x+1/x = 61/30 but i'm not too sure that's the right way

OpenStudy (phi):

that is the correct way. you can multiply both sides (all terms!) by x as the next step.

OpenStudy (mathstudent55):

Your equation is correct. Do you need help solving it?

OpenStudy (josefeth00):

If i multiply all the terms by x, i would get x^2 but what about the 1/x ?

OpenStudy (mathstudent55):

The x's will cancel there. \(\dfrac{1}{x} \times x = \dfrac{1}{x} \times \dfrac{x}{1} = \dfrac{x}{x} = 1\)

OpenStudy (josefeth00):

Okay I thought it would be that way I was just very confused. Thanks!

OpenStudy (phi):

don't forget to multiply the 61/30 by x also

OpenStudy (josefeth00):

Okay

OpenStudy (phi):

what do you have so far?

OpenStudy (josefeth00):

\[x ^{2} -\frac{ 61 }{ 30 }x +1 \]

OpenStudy (phi):

=0

OpenStudy (josefeth00):

I will be using quadratic formula to solve for x

OpenStudy (phi):

yes.

OpenStudy (josefeth00):

This question is on my textbook and it tells me I cannot use a calculator, is there any easy way to solve without having to multiply so many fractions?

OpenStudy (mathstudent55):

\(x \times x + \dfrac{1}{x} \times x = \dfrac{61}{30} \times x\) \(x^2 + 1 = \dfrac{61}{30} x\) \(x^2 - \dfrac{61}{30} x + 1 = 0\) Now continue with: \(x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (phi):

interesting. it almost sounds like there is a short-cut.

OpenStudy (josefeth00):

It is easy with a calculator but it is very weird without one

OpenStudy (phi):

we could multiply by 30 (both sides, all terms) to get \[ 30 x^2 -61x +30=0 \] but that still looks hard to do without a calculator.

OpenStudy (phi):

oh, we do get (relatively) nice numbers

OpenStudy (josefeth00):

I got it haha

OpenStudy (mathstudent55):